2. A mass M is connected by a massless rope going over a massless, frictionless pulley to a container of
water, which is initially at rest, and has mass m0. At time t = 0 the system is released and water is ejected
(with the help of internal mechanism) in the downward direction with a constant velocity U relative to
the container. The rate at which water is ejected is k m(t) where k is a constant. Find the acceleration of
mass M as a function of time (take M>m0).

Grade:12th pass

1 Answers

ASANKA
3 years ago
1*upward acting force on container due to loss of mass is = Ku
2*downward force acting on container due to mass at a time t =(m0-kt)g
3*downward acting force on mass M =Mg
4*net force= total mass at time t x acceleration (a)
so,  Mg+ku-(m0-kt)g=(M+m0-kt)a
thus we can find acceleration.